1-D Question 21 Solve the equation for exact solutions. arcsir(y - 3) } Posten . P67

2-Question 24 Find the phase shift of the function. y=-3-3 sin 2x units to the left units to the right units to the right units
Answer #1

1) Now We find the the cract value ar c sin ( y- 1) = *63 7- 1 = sin (1/3) - Y-7/6 V/ » Ya V} + 383th 6 Ya arc sin(y-7/6)= */